# Robust Bernoulli mixture models for credit portfolio risk

### v1

@@ Abstract
Keywords: Average Value-at-Risk, convex order, copula, factor model, positive dependence, tail dependence
1Department of Artificial Intelligence and Human Interfaces,
University of Salzburg Hellbrunner Strasse 34, 5020 Salzburg, Austria.
@@ 1 Introduction
A fundamental problem in modeling credit portfolio risk is that defaults are rare events, resulting in very limited available data. While the default probabilities of individual entities can generally be estimated with reasonable accuracy, e.g. using historical data or financial and macroeconomic indicators, and can therefore be treated as known, modeling the typically positive dependencies between default events poses a greater challenge. These dependencies, driven by common exposure to economic cycles, market shocks, and industry-specific risk factors, are complex to capture accurately due to the curse of dimensionality. For this reason, factor models for credit portfolio risk have been introduced, which are economically interpretable, easier to estimate, have well understood properties and are reliable under different scenarios. Well-known examples of industry credit risk models are CreditMetrics (cf. JPMorgan (1997)), KMV’s PortfolioManager (cf. Kealhofer and Bohn (2001)) and CreditRisk+ (cf. Credit Suisse Financial Products (1997)). The first two models belong to the class of Bernoulli mixture models where default is described by the event when the firm’s latent asset return falls below a given threshold reflecting its liabilities. In contrast, the CreditRisk+ model is an actuarial approach that approximates the Bernoulli default indicator with a Poisson random variable, resulting in a Poisson mixture model as an approximation of the Bernoulli mixture model. However, due to the scarcity of data, estimating the dependencies between defaults of different entities remains a challenging task, even in factor models. This leads to a certain amount of ambiguity in some parameter estimates such as the asset correlation coefficient. Additionally, asset returns often exhibit tail dependencies, which are typically underestimated by Gaussian factor models. Consequently, tail risks are likely to be underrepresented in portfolio losses. These challenges highlight the need to study robustness in credit risk models, accounting for both tail dependencies and model uncertainty.
@@
In this paper, we establish robustness results for Bernoulli mixture models, assuming that the conditional default probabilities of individual borrowers increase with a common risk factor, which may be either a latent random variable or an observable random variable, such as a market index. Our main contribution, Theorem 2.1, is a comparison result for credit portfolio losses where we provide straightforward and interpretable conditions on the conditional default probability functions that lead to a convex ordering of the losses. More specifically, the loss variable of portfolio $A$ is smaller in convex order than that of portfolio $B$ if the default events in portfolio $B$ exhibit a stronger positive dependence on the (potentially portfolio-specific) common risk factor than those in portfolio $A$. The proof relies on a general construction of stochastically increasing Bernoulli mixture models via copula-based threshold models. As an additional contribution of our paper, we provide a comparison of credit portfolio losses in convex order through a pointwise comparison of the copulas underlying the threshold models (Proposition 4.1). We then address model uncertainty by establishing lower and upper bounds for credit portfolio losses in convex order (Theorem 4.5) and for credit portfolio risks in terms of law-invariant, convex risk measures (Corollary 4.7). Our robustness results enable a simple comparison of different models, allowing risk measures to be ranked across various model specifications. While most of the paper assumes that default probabilities increase with the common risk factor, Theorem 4.3 presents a robustness result that holds even without this monotonicity assumption. We illustrate our results in two contexts: first, when individual parameters are ambiguous, and second, when there is uncertainty about the copula modeling the dependence between default events. In both cases, we calculate bounds for the Average Value-at-Risk (AVaR), a benchmark risk measure in the Basel III and Solvency II frameworks, using both simulated and real portfolios.
@@
Our paper relates to a large literature on credit risk models. For a general introduction to credit risk modeling, we refer to Bluhm et al. (2010) among many others. Latent factor models have been studied in Koyluoglu and Hickman (1998), Crouhy et al. (2007) and Gordy (2000), who also point out the equivalence between threshold models and mixture models for the special case of CreditMetrics and CreditRisk+. Frey et al. (2001) relate factor models to copula models. More specifically, they show that the distribution of the number of defaults in a portfolio strongly depends on the underlying copula. We consider copula-based threshold models and link them to the class of stochastically increasing Bernoulli mixture models.
@@
Frey et al. (2001) also point out that the Gaussian copula is unable to reflect tail dependence (compare also McNeil et al. (2015), Section 11). The shortcomings of the Gaussian copula model for the pricing of credit derivatives are also discussed in Donnelly and Embrechts (2010), highlighting in particular the inability of the Gaussian copula model to reflect default clustering and the associated tail dependence. Crook and Moreira (2011) empirically document that non-Gaussian copula families can better model dependence structures of credit portfolios. The impact of the choice of the copula model in stress testing applications has been empirically studied, for example, in Koziol et al. (2015) for 17 German banks. Their results indicate that the Gaussian copula is suitable when analysing the effect of stress scenarios characterized by extreme losses and high correlations while heavy-tailed copulas are more appropriate in less extreme adverse scenarios. Eberlein et al. (2008) point out that more flexible (non-normal) factor distributions or the application of alternative copula models allow for a more complex dependence structure which may exhibit tail dependence. Non-Gaussian copula models for the modelling of dependent defaults in intensity models have first been introduced in Schönbucher and Schubert (2001). We consider flexible dependence structures described by general stochastically increasing copulas, including as important examples the Gaussian and various Archimedean copula settings.
@@
We build on the aforementioned literature by, on the one hand, extending standard industry models to a robust setting that accommodates model uncertainty and tail dependence. On the other hand, our robustness results for Bernoulli mixture models may also be seen in the context of improved credit portfolio risk bounds under the structural assumption of conditional independence—an assumption that is both tractable and flexible, allowing for a broad spectrum of positive dependencies. In addition, our ordering results enable a ranking of portfolios losses in terms of their credit risk, which provides a transparent and objective basis for the assessment of portfolio risk levels.
The rest of the paper is organized as follows. Section 2 establishes our main result—a convex ordering of credit portfolio losses for stochastically increasing Bernoulli mixture models. Section 3 presents the construction of this class of models via copula-based threshold models while Section 4 derives lower and upper bounds for credit portfolio losses. Section 5 illustrates the evidence our results in both simulated and real data examples. All proofs are deferred to the appendix.
@@ 2 Setting and Main Result
We consider a portfolio of $N$ loans to borrowers $n=1,\ldots,N$, and denote by $\pi_{n}$ the (unconditional) default probability and by $D_{n}\colon\Omega\to\{0,1\}$ the default indicator of borrower $n$, where $D_{n}=1$ indicates default and $D_{n}=0$ no default111 We generally assume that the underlying probability space $(\Omega,\mathcal{A},P)$ is atomless and, hence, admits a continuous random variable. . Let $Z$ be a real-valued random variable, which will serve as a common risk factor. As default models, we consider a class of Bernoulli mixture models (BMMs) $\{(D_{1},\ldots,D_{N},Z)\}\,,$ where
@@
Denote by $p_{D_{n}}(z):=P(D_{n}=1|Z=z)$ the conditional default probability function. If, additionally to (I) and (II),
@@
we denote $(D_{1},\ldots,D_{N},Z)$ as stochastically increasing Bernoulli mixture model (siBMM). The individual default probabilities $\pi_{1},\ldots,\pi_{N}$ can often be estimated and thus assumed to be known, so Assumption (I) is reasonable. Since default events are typically not independent, but rather positively dependent, the dependence structure between $D_{1},\ldots,D_{N}$ is crucial for modeling portfolio risks. We impose conditional independence in (II) as standard structural assumption for modeling simple dependencies, where the common risk factor $Z$ may be an observable factor, e.g. a market index, or a latent random variable that follows an arbitrary distribution, e.g. a discrete or continuous distribution. Assumption (III) incorporates positive dependence into our model. It states that, for all borrowers $n=1\ldots,N$, the conditional default probabilities increase with increasing value $z$ of the common risk factor, which equivalently means that the default indicator variable $D_{n}$ is stochastically increasing in $Z,$ see Definition 3.2 and (Müller and Stoyan, 2002, Section 3.10). This assumption is typically not imposed in the definition of a Bernoulli mixture model, but is fulfilled by many models, see (McNeil et al., 2015, Section 11.2). In practice, the risk factor $Z$ may be interpreted as a market index which has a positive or negative influence on the default probabilities. For example, if the market deteriorates, default events may be more likely. In the latter case, $Z$ is a decreasing transformation of a market index.
Later, we provide a copula-based construction of Bernoulli mixture models that imply conditional default probabilities which are increasing in the systematic risk factor (compare Proposition 3.3), i.e. we construct stochastically increasing Bernoulli mixture models as specified above. The increasingness Assumption (III) implies that any siBMM $(D_{1},\ldots,D_{N},Z)$ satisfies
$P(D_{n}\geq x,Z\geq z)$ $=\int^{\infty}_{z}\mathbb{P}(D_{n}\geq x\mid Z=z^{\prime})\mathrm{\,d}\mathbb{P}^{Z}(z^{\prime})$
$\geq\int^{\infty}_{z}\mathbb{P}(D_{n}\geq x)\mathrm{\,d}\mathbb{P}^{Z}(z^{\prime})=\mathbb{P}(D_{n}\geq x)\mathbb{P}(Z\geq z)\,,$
for all $x,z\in\mathbb{R}\,,$ which yields, in particular, a positive Pearson correlation of $D_{n}$ and $Z\,.$ Further, due to conditional independence, any siBMM has a pointwise non-negative correlation matrix for $(D_{1},\ldots,D_{N})\,.$ The assumptions of a siBMM allow to model various cases of positive dependence for $(D_{1},\ldots,D_{n})\,,$ in particular, independence (where the function $p_{D_{n}}$ is constant for all $n$) and, if $F_{Z}$ is continuous, also comonotonocity [^2] A bivariate random vector $(X,Y)$ is said to be comonotonic if there exist increasing functions $f\colon\mathbb{R}\to\mathbb{R}$ and $g\colon\mathbb{R}\to\mathbb{R}$ and a random variable $U$ such that $X=f(U)$ and $Y=g(U)$ almost surely. In this regard, comonotonicity models perfect positive dependence. (where $p_{D_{n}}$ takes the form $p_{D_{n}}(z)=\mathds{1}_{\{z\geq F_{Z}^{-1}(1-\pi_{n})\}}$ for all $n$). The latter case means that default is a deterministic function of $Z\,.$
@@
Let $e_{n}\in(0,\infty)$ denote the (deterministic) exposure at default associated with borrower $n$ and $\delta_{n}\in(0,1]$ its (percentage) loss given default variable which is assumed to be a random variable that is independent of all other sources of randomness. Then the bank’s total loss is given by
@@
Figure 2.1: Left plot: Examples of transformed conditional default probabilities $p_{D_{n}}\circ F_{Z}^{-1}=P(D_{n}=1\mid Z=F_{Z}^{-1}(\cdot))$ which are by Assumption (III) increasing and satisfy the constraint $\int_{0}^{1}P(D_{n}=1\mid Z=F_{Z}^{-1}(t))\mathrm{\,d}t=\pi_{n}$ with $\pi_{n}$ chosen as $0.2\,.$ The solid line models independence of $D_{n}$ and $Z\,,$ while the dashed line models perfect positive dependence of $D_{n}$ on $Z\,.$ Right plot: Default integral functions $G_{D_{n},Z}$ defined in (2.2) as the antiderivatives of the respective transformed conditional default probabilities from the left plot, noting that each $G_{D_{n},Z}$ is increasing convex and passes through the points $(0,0)$ and $(0,\pi_{n})\,.$ Under Assumption (III), the independent and comonotonic dependence relations mark the extreme cases (solid and dashed) while, for all other dependencies, the default integral functions lie between these extremes.
As a main contribution, we provide simple and interpretable conditions on the conditional default probabilities $p_{D_{n}}$ to imply a comparison of credit portfolio losses in convex order, where, for non-negative random variables $X$ and $Y,$ the convex order $X\leq_{cx}Y$ is defined by $\mathbb{E}\varphi(X)\leq\mathbb{E}\varphi(Y)$ for all convex functions $\varphi\colon[0,\infty)\to\mathbb{R}\,,$ see (Shaked and Shantikumar, 2007, Section 3.A) for an overview of the convex order. More precisely, our conditions involve a pointwise comparison of the default integral functions
$G_{D_{n},Z}(s):=\int_{0}^{s}P(D_{n}=1\mid Z=F_{Z}^{-1}(t))\mathrm{\,d}t,=\int_{0}^{s}p_{D_{n}}(F_{Z}^{-1}(t))\mathrm{\,d}t,\quad s\in[0,1]\,,$ (2.2)
associated with $p_{D_{n}}$. Here, $F_{Z}^{-1}$ denotes the generalized inverse of the distribution function of $Z\,,$ i.e., $F_{Z}^{-1}(t):=\inf\{x\mid F_{Z}(x)\geq t\}$ for $t\in(0,1)\,.$ It is obvious that $G_{D_{n},Z}$ is a continuous increasing function mapping from $(0,1)$ to $[0,\pi_{n}]$ such that $G_{D_{n},Z}(0)=0$ and $G_{D_{n},Z}(1)=\pi_{n}\,.$ Further, under Assumption (III), $G_{D_{n},Z}$ is convex and lies between the extreme cases of independence and comonotonicity as discussed before, see Figure 2.1. The following theorem is our main result.
@@
Consider default models $(D_{1},\ldots,D_{N},Z)$ and $(D_{1}^{\prime},\ldots,D_{N}^{\prime},Z^{\prime})$ that satisfy Assumptions (I)–(III). Let $\delta_{1},\ldots,\delta_{N}$ be independent non-negative random variables that are also independent of the default models and let $e_{1},\ldots,e_{N}$ be positive constants. If $G_{D_{n},Z}(s)\geq G_{D_{n}^{\prime},Z^{\prime}}(s)$ for all $n$ and $s\,,$ then
@@
The proof of Theorem 2.1 is deferred to the appendix and relies on a representation of siBMMs through copula-based threshold models, which we provide in the next section. As a consequence of the above theorem, we establish robustness results for credit risk models and determine in Theorem 4.5 lower and upper bounds in convex order for classes of credit portfolios models.
@@
[…]
@@
    The pointwise comparison of default integral functions in Theorem 2.1 allows a simple interpretation: Smaller values of $G_{D_{n},Z}$ indicate more positive dependence between $D_{n}$ and $Z.$ More precisely, if $Z$ has the same distribution as $Z^{\prime}\,,$ then $G_{D_{n},Z}(s)\geq G_{D_{n}^{\prime},Z^{\prime}}(s)$ for all $s$ is equivalent to
    $P(D_{n}>x,Z>z)\leq P(D_{n}^{\prime}>x,Z^{\prime}>z)\quad\text{for all }x,z$ (2.4)
    or to $P(D_{n}=1\mid Z>z)\leq P(D_{n}^{\prime}=1\mid Z^{\prime}>z)$ for all $z,$ i.e., default is more likely in the model $(D_{n}^{\prime},Z^{\prime})$ when the risk factor exceeds fixed values. The relation in (2.4) is well studied in the literature and known as upper orthant order or more positive quadrant dependence order, see e.g. (Rüschendorf, 2013, Chapter 6) or Joe (1997). Since the supermodular order discussed in (a) implies the upper orthant order, see (Müller and Stoyan, 2002, Figure 3.1), we also obtain under the assumptions of Theorem 2.1 that joint default probabilities are ordered, i.e., $P(D_{n}=1,n\in M)\leq P(D_{n}^{\prime}=1,n\in M)$ for all $M\subseteq\{1,\ldots,N\}\,.$
@@
    The focus of this paper is on Bernoulli mixture models under Assumption (III) stating that default probabilities are stochastically increasing in the common risk factor. However, if this assumption is dropped, the credit portfolio loss of a BMM is still dominated by the credit portfolio loss of a siBMM when considering the increasing rearrangement of the conditional default probability function $p_{D_{n}}$. This is the content of Theorem 4.3.
  4. (d)

### v2

@@ Abstract
Keywords: Average Value-at-Risk, convex order. copula, CreditMetrics, factor model, KVM Portfolio Manager, model uncertainty, positive dependence, tail dependence
1Department of Mathematics,
University of Salzburg, Hellbrunner Strasse 34, 5020 Salzburg, Austria.
@@ 1 Introduction
A fundamental problem in modeling credit portfolio risk is that defaults are rare events, resulting in very limited available data. While the default probabilities of individual entities can generally be estimated with reasonable accuracy, e.g. using historical data or financial and macroeconomic indicators, and can therefore be treated as known, modeling the typically positive dependencies between default events poses a greater challenge. These dependencies, driven by common exposure to economic cycles, market shocks, and industry-specific risk factors, are complex to capture accurately due to the curse of dimensionality. For this reason, factor models for credit portfolio risk have been introduced, which are economically interpretable, easier to estimate, have well understood properties, and are reliable under different scenarios. Well known examples of industry credit risk models are CreditMetrics (cf. JPMorgan (1997)), KMV’s PortfolioManager (cf. Kealhofer and Bohn (2001)), and CreditRisk+ (cf. Credit Suisse Financial Products (1997)). The first two models belong to the class of Bernoulli mixture models (BMMs) where default is described by the event when the firm’s latent asset return falls below a given threshold reflecting its liabilities. In contrast, the CreditRisk+ model is an actuarial approach that approximates the Bernoulli default indicator with a Poisson random variable, resulting in a Poisson mixture model as an approximation of the BMM. However, due to the scarcity of data, estimating the dependencies between defaults of different entities remains a challenging task, even in factor models. This leads to ambiguity in some parameter estimates such as the asset correlation coefficient. Additionally, asset returns often exhibit tail dependencies, which are typically underestimated by the widely used Gaussian factor models. Consequently, tail risks are likely to be underrepresented in simulated portfolio losses. These challenges highlight the need to study robustness in credit risk models, accounting for both tail dependencies and model uncertainty.
@@
In this paper, we establish robustness results for stochastically increasing Bernoulli mixture models (siBMMs), where the conditional default probabilities of individual borrowers increase with a common risk factor, which may be either a latent random variable or an observable random variable, such as a market index. Our main contribution, Theorem 2.1, is a comparison result for credit portfolio losses where we provide straightforward and interpretable conditions on the conditional default probability functions that lead to a convex ordering of the losses. More specifically, the loss variable of portfolio $A$ is smaller in convex order than that of portfolio $B$ if the default events in portfolio $B$ exhibit a stronger positive dependence on the (potentially portfolio-specific) common risk factor than those in portfolio $A$. The proof relies on a general construction of siBMMs via copula-based threshold models. While most of the paper assumes that default probabilities increase with the common risk factor, Theorem 2.3 presents a robustness result that holds even without this monotonicity assumption. We then address model uncertainty by establishing lower and upper bounds for credit portfolio losses in convex order (Theorem 3.1) and for credit portfolio risks in terms of law-invariant, convex risk measures (Corollary 3.3). Our robustness results enable a simple comparison of different models, allowing risk measures to be ranked across various model specifications. As an additional contribution of our paper, we provide a comparison of credit portfolio losses in convex order through a pointwise comparison of the copulas underlying the threshold models (Proposition A.6). We illustrate our results in two contexts: first, when individual parameters are ambiguous, and second, when there is uncertainty about the copula modeling the dependence between default events. In both cases, we calculate bounds for the Average Value-at-Risk (AVaR), a benchmark risk measure in the Basel III and Solvency II frameworks, using both simulated and real portfolios.
@@
Our paper relates to a large literature on credit risk models. For a general introduction to credit risk modeling, we refer to Bluhm et al. (2010) among many others. Latent factor models have been studied in Koyluoglu and Hickman (1998), Crouhy et al. (2000), and Gordy (2000), who also point out the equivalence between threshold models and mixture models for the special case of CreditMetrics and CreditRisk+. Frey et al. (2001) relate factor models to copula models. More specifically, they show that the distribution of the number of defaults in a portfolio strongly depends on the underlying copula. We consider copula-based threshold models and link them to the class of siBMMs.
@@
Frey et al. (2001) also point out that the Gaussian copula is unable to reflect tail dependence (compare also McNeil et al. (2015), Section 11). Donnelly and Embrechts (2010) highlight the inability of the Gaussian copula model to reflect default clustering and the associated tail dependence. Crook and Moreira (2011) empirically document that non-Gaussian copula families can better model dependence structures of credit portfolios. The impact of the choice of the copula model in stress testing applications has been empirically studied, for example, in Koziol et al. (2015). Their results indicate that the Gaussian copula is suitable when analysing the effect of stress scenarios characterized by extreme losses and high correlations while heavy-tailed copulas are more appropriate in less extreme adverse scenarios. Non-Gaussian copula models for the modelling of dependent defaults in intensity models have first been introduced in Schönbucher and Schubert (2001). We consider flexible dependence structures described by general stochastically increasing copulas, including as important examples the Gaussian and various Archimedean copula settings.
@@
We build on the aforementioned literature by, on the one hand, extending standard industry models to a robust setting that accommodates model uncertainty and tail dependence. On the other hand, our robustness results for BMMs may also be seen in the context of improved credit portfolio risk bounds under the structural assumption of conditional independence—an assumption that is both tractable and flexible, allowing for a broad spectrum of positive dependencies. In addition, our ordering results enable a ranking of portfolio losses in terms of their credit risk, which provides a transparent and objective basis for the assessment of portfolio risk levels.
The rest of the paper is organized as follows. Section 2 establishes our main result—a convex ordering of credit portfolio losses for siBMMs. Section 3 derives lower and upper bounds for credit portfolio losses. Section 4 illustrates our results in both simulated and real data examples. Appendix A provides a representation of siBMMs via copula-based threshold models, while all proofs are deferred to Appendix B.
@@ 2 Setting and Main Result
We consider a portfolio of $N$ loans to borrowers $n=1,\ldots,N$, and denote by $\pi_{n}$ the (unconditional) default probability and by $D_{n}\colon\Omega\to\{0,1\}$ the default indicator of borrower $n$, where $D_{n}=1$ indicates default and $D_{n}=0$ no default [^1] We generally assume that the underlying probability space $(\Omega,\mathcal{A},P)$ is atomless and, hence, admits a continuous random variable. . Let $Z$ be a real-valued random variable, which will serve as a common risk factor. As default models, we consider a class of Bernoulli mixture models (BMMs) $\{(D_{1},\ldots,D_{N},Z)\}\,$, where
@@
Denote by $p_{D_{n}}(z):=\mathbb{P}(D_{n}=1\mid Z=z)$ the conditional default probability function. If, additionally to (I) and (II),
@@
we refer to $(D_{1},\ldots,D_{N},Z)$ as stochastically increasing Bernoulli mixture model (siBMM), noting that a random variable $X$ is stochastically increasing (SI) in $Y$ if, for all $x\in\mathbb{R}\,$, $\mathbb{P}(X\geq x\mid Y=y)$ is increasing in $y$. The individual default probabilities $\pi_{1},\ldots,\pi_{N}$ can often be estimated and thus assumed to be known, so Assumption (I) is reasonable. Since default events are typically not independent, but rather positively dependent, the dependence structure between $D_{1},\ldots,D_{N}$ is crucial for modeling portfolio risks. We impose conditional independence in (II) as standard structural assumption for modeling simple dependencies, where the common risk factor $Z$ may be an observable factor, e.g. a market index, or a latent random variable that follows an arbitrary distribution, e.g. a discrete or continuous distribution. Assumption (III) incorporates positive dependence into our model, stating that the default indicators $D_{n}$ are stochastically increasing in $Z.$ This assumption is typically not imposed in the definition of a BMM, but is fulfilled by many models; see (McNeil et al., 2015, Section 11.2). In practice, the risk factor $Z$ may be interpreted as a market index which has a positive or negative influence on the default probabilities. For example, if the market deteriorates, default events may be more likely. In the latter case, $Z$ is a decreasing transformation of a market index. The assumptions of a siBMM allow to model various cases of positive dependence for $(D_{1},\ldots,D_{n})\,$, in particular, independence (where the function $p_{D_{n}}$ is constant for all $n$) and, if $F_{Z}$ is continuous, also comonotonocity [^2] A bivariate random vector $(X,Y)$ is said to be comonotonic if there exist increasing functions $f\colon\mathbb{R}\to\mathbb{R}$ and $g\colon\mathbb{R}\to\mathbb{R}$ and a random variable $U$ such that $X=f(U)$ and $Y=g(U)$ almost surely. In this regard, comonotonicity models perfect positive dependence. (where $p_{D_{n}}$ takes the form $p_{D_{n}}(z)=\mathds{1}_{\{z\geq F_{Z}^{-1}(1-\pi_{n})\}}$ for all $n$). The latter case means that default is a deterministic function of $Z.$
@@
Let $e_{n}\in(0,\infty)$ denote the (deterministic) exposure at default of borrower $n$ and $\delta_{n}\in(0,1]$ its (percentage) loss given default, which is assumed to be a random variable that is independent of all other sources of randomness. Then the bank’s total loss is given by
@@
Figure 2.1: Left plot: Some examples of transformed conditional default probabilities $p_{D_{n}}\circ F_{Z}^{-1}=\mathbb{P}(D_{n}=1\mid Z=F_{Z}^{-1}(\cdot))$ which are by Assumption (III) increasing and satisfy the constraint $\int_{0}^{1}\mathbb{P}(D_{n}=1\mid Z=F_{Z}^{-1}(t))\mathrm{\,d}t=\pi_{n}$ with $\pi_{n}$ chosen as $0.2.$ The solid line models independence of $D_{n}$ and Z,Z\,, while the dashed line models perfect positive dependence of $D_{n}$ on $Z.$ Right plot: Default integral functions $G_{D_{n},Z}$ defined in (2.2) as the antiderivatives of the respective transformed conditional default probabilities from the left plot, noting that each $G_{D_{n},Z}$ is increasing convex and passes through the points $(0,0)$ and $(0,\pi_{n}).$ Under Assumption (III), the independent and comonotonic dependence relations mark the extreme cases (solid and dashed) while, for all other dependencies, the default integral functions lie between these extremes.
As a main contribution, we provide simple and interpretable conditions on the conditional default probabilities $p_{D_{n}}$ to imply a comparison of credit portfolio losses in convex order. Recall that, for non-negative random variables $X$ and $Y,$ the convex order $X\leq_{cx}Y$ is defined by $\mathbb{E}\varphi(X)\leq\mathbb{E}\varphi(Y)$ for all convex functions $\varphi\colon[0,\infty)\to\mathbb{R};$ see (Shaked and Shantikumar, 2007, Section 3.A) for an overview of the convex order. More precisely, our conditions involve a pointwise comparison of the default integral functions
$G_{D_{n},Z}(s):=\int_{0}^{s}\mathbb{P}(D_{n}=1\mid Z=F_{Z}^{-1}(t))\mathrm{\,d}t=\int_{0}^{s}p_{D_{n}}(F_{Z}^{-1}(t))\mathrm{\,d}t,\quad s\in[0,1],$ (2.2)
associated with $p_{D_{n}}$. Here, $F_{Z}^{-1}$ denotes the generalized inverse of the distribution function of Z,Z\,, i.e., $F_{Z}^{-1}(t):=\inf\{x\mid F_{Z}(x)\geq t\}$ for $t\in(0,1).$ Thus, $G_{D_{n},Z}(s)$ gives the average conditional default probability over the lowest $s$-quantile of the risk factor $Z$. It is obvious that $G_{D_{n},Z}$ is a continuous increasing function mapping from $(0,1)$ to $0,\pi_{n}]$ such that $G_{D_{n},Z}(0)=0$ and $G_{D_{n},Z}(1)=\pi_{n}.$ Further, under Assumption (III), $G_{D_{n},Z}$ is convex and lies between the extreme cases of independence and comonotonicity as discussed before; see Figure 2.1. The following theorem is our main result.
@@
Consider default models $(D_{1},\ldots,D_{N},Z)$ and $(D_{1}^{\prime},\ldots,D_{N}^{\prime},Z^{\prime})$ that satisfy Assumptions (I)–(III) with unconditional default probabilities $(\pi_{1},\ldots,\pi_{N})$. Let $\delta_{1},\ldots,\delta_{N}$ be independent non-negative random variables that are also independent of the default models and let $e_{1},\ldots,e_{N}$ be positive constants. If $G_{D_{n},Z}(s)\geq G_{D_{n}^{\prime},Z^{\prime}}(s)$ for all $n$ and $s,$ then
@@
As a consequence of the above theorem, we establish in Section 3 robustness results for credit risk models and determine in Theorem 3.1 lower and upper bounds in convex order for classes of credit portfolios models.
@@
    The pointwise comparison of default integral functions in Theorem 2.1 allows a simple interpretation: Smaller values of $G_{D_{n},Z}$ indicate more positive dependence between $D_{n}$ and $Z.$ More precisely, if $Z$ has the same distribution as $Z^{\prime},$ then $G_{D_{n},Z}(s)\geq G_{D_{n}^{\prime},Z^{\prime}}(s)$ for all $s$ is equivalent to $\mathbb{P}(D_{n}>x,Z>z)\leq\mathbb{P}(D_{n}^{\prime}>x,Z^{\prime}>z)$ for all $x,z$ or to $\mathbb{P}(D_{n}=1\mid Z>z)\leq\mathbb{P}(D_{n}^{\prime}=1\mid Z^{\prime}>z)$ for all $z,$ i.e., default is more likely in the model $(D_{n}^{\prime},Z^{\prime})$ when the risk factor exceeds fixed values. The above relation is well studied in the literature and known as upper orthant order or more positive quadrant dependence order; see e.g. (Rüschendorf, 2013, Chapter 6) or Joe (1997).
@@
    The proof of Theorem 2.1 is based on a representation of siBBMs through threshold models (see Appendix A) and on a supermodular comparison result for factor models (see (Ansari and Rüschendorf, 2024, Corollary 4.1)). Since the supermodular order implies the upper orthant order, see (Müller and Stoyan, 2002, Figure 3.1), we also obtain under the assumptions of Theorem 2.1 that joint default probabilities are ordered, i.e., $\mathbb{P}(D_{n}=1,n\in M)\leq\mathbb{P}(D_{n}^{\prime}=1,n\in M)$ for all $M\subseteq\{1,\ldots,N\}.$